Circular No. 8526

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 EL_{173} (LONEOS) An apparently asteroidal object that was discovered by the LONEOS

project, and designated 2005 EL_{173} by the Minor Planet Center (discovery observation below from MPS 129296), has been found by A. Fitzsimmons (Queen’s University, Belfast) to have an asymmetric coma extending to 3′′.5 in p.a. 70^{o} on combined CCD R-band frames (total exposure 1650 s) taken on May 10.0 UT with the 3.6-m New Technology Telescope at the European Southern Observatory; individual sub-arcsecond 110-s frames show a compact coma.

2005 UT  α_{2000}  δ_{2000}  Mag.
Mar. 8.24106  11^{h}14^{m}23.^{s}44  +7^{o}40′49′′.6  19.3

The following hyperbolic orbital elements by B. G. Marsden, Smithsonian Astrophysical Observatory, are from 39 observations spanning 2005 Mar. 3–May 4 (including prediscovery Spacewatch observations from Mar. 3; mean residual 0′′.5). The “original” barycentric value of 1/a is +0.000113 ± 0.000056 AU^{−1} (1σ mean error), while the “future” value of 1/a is +0.000031 AU^{−1}.

Epoch = 2007 Mar. 1.0 TT

T = 2007 Mar. 6.1192 TT  ω = 261.^{o}3354
e = 1.003109  Ω = 344.7938 }2000.0
q = 3.897440 AU  i = 130.7181
2005 EO_{304}

S. D. Kern and J. L. Elliot, Massachusetts Institute of Technology, report that V R-band observations, obtained with the 6.5-m Clay telescope (+ MagIC) on Apr. 15 UT in ∼ 0′′.7 seeing, reveal 2005 EO_{304} (cf. MPECs 2005-G51, 2005-J15; MPS 130312) to be a binary system. These observations are part of an ongoing program for recovery of new tranneptunian objects discovered by the ‘Deep Ecliptic Survey’. The fainter member of the pair lies 2′′.67 ± 0′′.06 from the brighter member at p.a. 105^{o} ± 1^{o}. The primary member of the pair is 1.2 ± 0.1 magnitudes brighter than its companion. Additional photometric and astrometric observations of 2005 EO_{304} are strongly encouraged.

2005 May 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green